OpenStudy (anonymous):

Find the 6th partial sum of a 78,432 b 2,387 c 9,555 d 11,204

OpenStudy (anonymous):

\[\sum_{i=1}^{\infty} 7(4)^{i-1}\]

OpenStudy (amistre64):

factor out the constants

OpenStudy (amistre64):

7/4 (4^1+4^2+...) which is a geometric sum if anything right?

OpenStudy (anonymous):

An \(n\)th partial sum is the evaluation of the sum to \(n\) terms. So, evaluate \[\sum_{1 \leq i \leq 6}7(4)^{i-1}\]

OpenStudy (amistre64):

if we go ahead and factor out another 4 to make the series start at 4^0 we get: 7(4^0 + ... 4^5) i believe

OpenStudy (anonymous):

I have no idea what you are saying

OpenStudy (amistre64):

\[7(4^0)+7(4^1)+7(4^2)+7(4^3)+7(4^4)+7(4^5)\] \[7(1+4+16+...+4^5)\]

OpenStudy (anonymous):

Okay amistre64 why am I doing this? I know what you are doing, just not why

OpenStudy (amistre64):

im trying to simplify it out to get to the geometric part that i can use the formula on

OpenStudy (anonymous):

Okay

OpenStudy (amistre64):

hence:\[7*\frac{1-4^5}{1-4}\] if i recall it right

OpenStudy (anonymous):

Okay, then what?

OpenStudy (amistre64):

or is it 4^6 up there? hmm

OpenStudy (anonymous):

It should be 4^5 because i am looking for the sixth. right

OpenStudy (amistre64):

1 4 4^2 ... 4^5 -(4 4^2 ... 4^5 4^6) im going with 4^6

OpenStudy (amistre64):

16*16*16 need a calculator lol

OpenStudy (anonymous):

That is 4096

OpenStudy (anonymous):

\[ \sum_{1 \leq k \leq n}ar^k=a\frac{1-r^{n}}{1-r} \] \[ \sum_{1 \leq i \leq 6}7(4)^{i-1}=\sum_{0 \leq i \leq 5}7(4)^i=7\frac{1-4^5}{1-4} \]

OpenStudy (anonymous):

AHHHH! I am so confused!!

OpenStudy (amistre64):

-1 and divide by 3 then

OpenStudy (amistre64):

i get 9555 over all

OpenStudy (anonymous):

Limitless, what did you get?

OpenStudy (anonymous):

Okay, because Limitless is gone now, I guess I am going with 9555

OpenStudy (anonymous):

Apparently the geometric sum formula is wrong. It should be \[ \sum_{0 \leq i \leq n}ar^i=\frac{a(k^{n+1}-1)}{k-1} \] So: \[ \sum_{0 \leq i \leq 5}7(4)^i=\frac{7(4^{5+1}-1)}{4-1}=\frac{7(4^{6}-1)}{3}=9,555\]

OpenStudy (anonymous):

Okay, thanks guys!